I am having a hard time with this. I don't understand. Can someone please break down and help me answer the below questions? I have viewed lectures and looked at books but I am getting no where.
2. Given the line with the equation y= -2, Find three ordered pairs to satisfy the equation. Find the slope of the line. Graph the line.
3. Given the line with equation x =3, Find 3 ordered pairs that satisfy the equation. Find slope of line. Graph the line.
4. Given the line with equation 2x + 5y =10
write equation in slope-intercept form. state slope of the line. state y-intercept of line as ordered pair. Use slope-intercept form to graph line.
5. Find slope of line through points (-3, 7) and (2, -5)
6. Write in point slope form, y - y^1 = m (x - x^1), equation of line through the two points given in above question.
7. Convert equation that wrote in above question to slope intercept form, y = mx + b
8. convert equation in #6 to standard form, Ax + By = C, A, B, C are integers and A > 0
9. Given line with equation 3x-y=5, write equation in point-slope form of line perpendicular to given line through point )3, -2)
10. ) The manager at a local Best Buy store kept track of daily sales of the iPhone 6 when it was first introduced. He noticed that sales were very high at first, but showed a steady decrease. For example, on day 5 after the phones first went on sale, the store sold 238 iPhone 6’s, but on day 10 after the phones first went on sale, the store sold only 153 iPhone 6’s. Assume that the relationship between number of days since the phone became available and number sold is linear and, let x = the number of days since the phone was released and y = number of phones sold at this Best Buy store.

a) Write two ordered pairs to represent the sales on the days described.
b) Find the slope of the line through those ordered pairs.
c) Write an equation in point-slope form of the line.
d) Convert the equation to slope-intercept form.
e) Explain what the slope and the y-intercept mean in the context of this problem.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

2. Given the line with the equation y= -2, Find three ordered pairs to satisfy the equation. Find the slope of the line. Graph the line.
---
y = -2 for all values of x.
eg, (0,-2), (-6,-2), (100,-2)
It's parallel to the x-axis --> slope = 0
==============
3. Given the line with equation x =3, Find 3 ordered pairs that satisfy the equation. Find slope of line. Graph the line.
Similar to #2.
Parallel to the y-axis --> slope is infinite, or undefined.
----------------
4. Given the line with equation 2x + 5y =10
write equation in slope-intercept form. state slope of the line. state y-intercept of line as ordered pair. Use slope-intercept form to graph line.
---
Slope-intercept means solve for y.
2x + 5y = 10
5y = -2x + 10
y = (-2/5)x + 2
Slope is the coefficient of x = -2/5
y = mx + b, where m = slope and b is the y-intercept
y-int = 2 or (0,2)
===============================
5. Find slope of line through points (-3,7) and (2,-5)
Slope m = diffy/diffx
m = (7 - (-5))/(-3-2) = 12/-5
m = -12/5
======================
6. Write in point slope form, y - y^1 = m (x - x^1), equation of line through the two points given in above question.
^ is used for exponents.
y - y1 = m (x - x1)
Use either point.
y - 7 = (-12/5)*(x + 3)
=========================
7. Convert equation that wrote in above question to slope intercept form, y = mx + b
----
y - 7 = (-12/5)*(x + 3)
y-7 = (-12/5)*x - 36/5
y = (-12/5)x - 1/5
=====================
8. convert equation in #6 to standard form, Ax + By = C, A, B, C are integers and A > 0
-----
y = (-12/5)x - 1/5
5y = -12x - 1
12x + 5y = -1
===================
9. Given line with equation 3x-y=5, write equation in point-slope form of line perpendicular to given line through point )3, -2)
---
Step 1, find the slope m.
The slope of all lines perpendicular is the negative inverse of m.
------
Step 2, same as #6 above.
==============================
10. ) The manager at a local Best Buy store kept track of daily sales of the iPhone 6 when it was first introduced. He noticed that sales were very high at first, but showed a steady decrease. For example, on day 5 after the phones first went on sale, the store sold 238 iPhone 6’s, but on day 10 after the phones first went on sale, the store sold only 153 iPhone 6’s. Assume that the relationship between number of days since the phone became available and number sold is linear and, let x = the number of days since the phone was released and y = number of phones sold at this Best Buy store.
a) Write two ordered pairs to represent the sales on the days described.
b) Find the slope of the line through those ordered pairs.
c) Write an equation in point-slope form of the line.
d) Convert the equation to slope-intercept form.
e) Explain what the slope and the y-intercept mean in the context of this problem.

Answer by stanbon(75887)

2. Given the line with the equation y= -2, Find three ordered pairs to satisfy the equation. Find the slope of the line. Graph the line.
Plot the points:: (0,-2), (2,-2) and draw a line thru them
The equation is y = 0*x-2, so the slope is m = 0
----------------------------------------------------
3. Given the line with equation x =3, Find 3 ordered pairs that satisfy the equation. Find slope of line. Graph the line.
Plot the points:: (3,0) and (3,5) and draw vertical line thru them.
Since the line is vertical, the slope is undefined.
----------------------------------------
4. Given the line with equation 2x + 5y =10
write equation in slope-intercept form. state slope of the line. state y-intercept of line as ordered pair. Use slope-intercept form to graph line.
5y = -2x+10
y = (-2/5)x + 2
Plot the points:: (0,2), (5,0) and draw a line thru them.
---------------------------
5. Find slope of line through points (-3, 7) and (2, -5)
m = (-5-7)/(2--3) = -12/5 = -2.4
---------------------------------------------
6. Write in point slope form, y - y1 = m (x - x1), equation of line through the two points given in above question.
y-7 = -2.4(x--3)
y -7 = -2.4(x+3)
y = -2.4x - 14.2
--------------------------
7. Convert equation that wrote in above question to slope
intercept form, y = mx + b
Done
8. convert equation in #6 to standard form, Ax + By = C, A, B, C are integers and A > 0
(12/5)x + y = -71/5
12x + 5y = -71
-----------------
9. Given line with equation 3x-y=5, write equation in point-slope form of line perpendicular to given line through point (3, -2)
y = 3x-5
slope = 3
----
A perpendicular line must have slope = -1/3
Form:: y = mx + b
Solve for "b"
-2 = (-1/3)3 + b
b = -1
Ans: y = (-1/3)x -1
---------------------

10. ) The manager at a local Best Buy store kept track of daily sales of the iPhone 6 when it was first introduced. He noticed that sales were very high at first, but showed a steady decrease. For example, on day 5 after the phones first went on sale, the store sold 238 iPhone 6’s, but on day 10 after the phones first went on sale, the store sold only 153 iPhone 6’s.
Points:: (5,238), (10,153)
--------------------------
Assume that the relationship between number of days since the phone became available and number sold is linear and, let x = the number of days since the phone was released and y = number of phones sold at this Best Buy store.
a) Write two ordered pairs to represent the sales on the days described.
done above
----------------
I'll leave the rest of this to you.

b) Find the slope of the line through those ordered pairs.
c) Write an equation in point-slope form of the line.
d) Convert the equation to slope-intercept form.
e) Explain what the slope and the y-intercept mean in the context of this problem.
Ans: x-int::: # of the day when sales will be zero
y-int::: # of sales on day zero
-----
Cheers,
Stan H.
-------------